OpenStudy (anonymous):

What value completes the square for the expression? x^2 - 18x A. 9 B. -9 C. 81 D. -81

OpenStudy (anonymous):

@tgawade that other one is too messy

OpenStudy (anonymous):

add \[\left( \frac{ co-efficient ~of ~x }{ 2 } \right)^2\]

OpenStudy (anonymous):

Okay use this formula Third term = (middle term(-18) / 2)^2 So we get third term = 81 Now we have third term = 81 so just add & subtract it from equation x^2 - 18x + 81 - 81 so we now factor out x^2 - 18x + 18 =(x-9)^2 there fore we get (x-9)^2 - 81

OpenStudy (anonymous):

so the answer is b ?

OpenStudy (anonymous):

yes ? no ?

OpenStudy (anonymous):

if we add 81 then it get complete

OpenStudy (anonymous):

that doesn't answer my question

OpenStudy (anonymous):

so are you saying the answer is c then ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wow that took a long time